[Hallc_running] [New Logentry] RC update for Hall C, Friday swing to Saturday swing

mack at jlab.org mack at jlab.org
Sat Mar 7 18:25:02 EST 2020

Logentry Text:
     Once we got going again Friday after accelerator issues and a Moeller measurement, took HMS/SHMS -3.5/-3.4 GeV data with 3He target, longitudinal polarization. Good production for remainder of Fri swing and Saturday owl. 

    Saturday day shift, the SHMS dipole PS started tripping repeatedly. Continued taking data with HMS of course, but beam trippy and less than 30 muA. 

    Switched to transverse target polarization and took production data with HMS only. In parallel, Steve L tried a software fix for the SHMS dipole PS, but no joy. 

    We were at the point of giving up on the SHMS for the weekend, when Jack S offered to bypass a relay in the PS. He was in the hall only 40min and the SHMS magnets were then set with no problems. https://logbooks.jlab.org/entry/3799371 

    The target polarization was switched back to longitudinal at the end of day shift. There was some confusion about this in the pass-down to the next shift, but the hclog documentation has been cleaned up. We're ba-a-a-ckkkkk in longitudinal production HMS/SHMS -3.5/-3.4 GeV data with 3He target.  Thanks to Steve L, Eric S, and Jack for spending a chunk of an otherwise very fine Saturday on the SHMS power supply issue. 

    Four target polarization measurements since restoration yesterday: 55.6%, 55%, 55.1%, and 54.2% (this last being transverse).


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3799443
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200307/ccb45444/attachment.html>

More information about the Hallc_running mailing list